Use the following information to answer questions 1 - 3

Exhale, Inc.
2012 Income Statement

  Net sales

$

9,200

  Cost of goods sold

7,600  

  Depreciation

350

  Earnings before interest and taxes

$

1,250  

  Interest paid

    35

  Taxable Income

$

1,215  

  Taxes

480  

  Net income

$

   735  

     Dividends

$

195

Exhale, Inc.
2012 Balance Sheet

2012

2012

  Cash

$

3,800    

  Accounts payable

$

3,420    

  Accounts rec.

1,100    

  Long-term debt

350    

  Inventory

4,100    

  Common stock

$

4,200    

  Total

$

9,000    

  Retained earnings

5,830    

  Net fixed assets

4,800    

  Total assets

$

13,800    

  Total liabilities & equity

$

13,800

Exhale, Inc., is currently operating at maximum capacity. All costs, assets, and current liabilities vary directly with sales. The tax rate and the dividend payout ratio will remain constant. In 2013, no new equity will be raised and sales are projected to increase by 10 percent. Construct the pro formas for 2013 (at first leave interest and long term debt unchanged). Then answer the following questions.

Projected total assets = $______

Hint: Divide each quantity on the left-hand side of the 2012 Balance Sheet by the 2012 Sales to compute the percentages of sales. New sales in 2013 will be 10% higher than in 2012, so compute sales for 2013. Apply the percentages you computed for 2012 to the new sales number for 2013 to construct each line item on the Balance Sheet for 2013. Read off the Total Assets at the bottom.

Projected 2013 Retained Earnings = $______

Hint. Proceed by constructing the Income Statement for 2013. Start with the new Sales (10% higher than 2012). Compute Cost of Goods Sold and Depreciation as percentages of sales in 2012 and apply the same percentages to 2013 Sales to get CoGS and Depr for 2013. When you get to taxes, things will be different. Compute the Taxes for 2012 as % of Taxable Income, not Sales, for 2012. Apply the same percentage to the 2013 Taxable Income to get Taxes for 2012. Similarly, compute Dividends for 2012 as % of 2012 Net Income, not Sales. Apply that percentage to the 2013 Net Income to get 2013 Dividends. The Addition to the Retained Earnings for 2013 is equal to the Net Income minus Dividends. Add the Addition to RE to the 2012 Retained Earnings from the Balance Sheet and you have the new 2013 Retained Earnings.

Additional new debt required = $______

Hint: You will need the answers of Q1 and Q2, but first you need to construct the right-hand side of the 2013 Balance Sheet. First, Accounts Payable as % of sales, so apply the 2012 % of sales for A/P to the new 2013 sales. Second, leave the LT Debt unchanged. Third, leave the Common Stock unchanged.Fourth, write in the new Retained Earnings from Q3. Compute the Total right-hand side of the B/S. It does not equal to the Total Assets from Q1. The difference is the additional debt to make the B/S balance.
Answer #1
INCOME STATEMENT
2012 Proforma2013
Net Sales $9,200 $10,120 (9200*1.1)
Cost of goods sold $7,600 $8,360 (7600*1.1)
Depreciation $350 $385 (350*1.1)
Earning Before interest and taxes $1,250 $1,375
Interest paid $35 $35
(480/1215)=39.5% Taxable income $1,215 $1,340
Taxes $480 $529 (1340*39.5%)
Net Income $735 $811
(195/735)=26.53% Dividend $195 $215 (811*26.53%)
Increase in retained Earnings $596 (811-215)
BALANCE SHEET A B=A/9200 C=B*10120 D E=D/9200
2012 Percentage of Sales Proforma2013 2012 Percentage of Sales Proforma2013
Cash $3,800 41.30% $4,180 Accounts Payable $3,420 37.17% $3,762 (10120*37.17%)
Accounts Receivable $1,100 11.96% $1,210 Long Term Debt $350 $792 (15180-10626-3762)
Inventory $4,100 44.57% $4,510 Common Stock $4,200 $4,200
Total $9,000 97.83% $9,900 Retained Earnings $5,830 $6,426 (5830+596))
Net Fixed assets $4,800 52.17% $5,280 Shareholders Equity $10,030 $10,626
Total Assets $13,800 150.00% $15,180 Total Liabilites& Equity $13,800 $15,180
Projected Total Assets $15,180
Projected 2013 Retained Earnings $6,426
Additional New Debt Required $442 (792-350)
Additional New Debt Required also can be calculated as shown below:
Increase in total assets $1,380 (15180-13800)
Less: Increase in retained Earnings $596
Less: Increase in Accounts Payable $342 (3762-3420)
Additional New Debt Required $442